OpenStudy (anonymous):

What is the standard form equation of the line passing through (–2, 8) with a slope of 2?

OpenStudy (anonymous):

y=mx+b.........y=2x-b..........plug in (x,y) to solve b ......such that 8=2(-2)+b ......b=12

OpenStudy (anonymous):

y=2x+12

OpenStudy (anonymous):

y = 2x+12 y = mx+b whre m is the slope.

OpenStudy (anonymous):

A. 2x + y = 12 B. 2x – y = 12 C. 2x + y = –12 D. 2x – y = –12 These are my choices..

OpenStudy (phi):

http://en.wikipedia.org/wiki/Linear_equation Standard form is Ax +By =C

OpenStudy (phi):

in y=2x+12 move the 2x to the left hand side

OpenStudy (anonymous):

But the 2x can't be negative.

OpenStudy (phi):

then multiply both sides by -1.

OpenStudy (anonymous):

So then the answer would be D?

OpenStudy (phi):

yes. y= 2x+12 -2x + y= 12 2x -y = -12